An object has a weight of 8N in air. However, it apparently weighs only 4N when it is completely submerged in water. What is the density of the object?
Answer #1

buoyant force

Fb = 8 - 4 = 4 N

1000* V * 9.8 = 4

V = 0.000408 m^3

as we also know that

8 - 4 = V g ( sigma - rho)

4 = 0.000408* 9.8* ( sigma - 1000)

sigma = 2000.4 kg/m^3

========